### v2

# Ionization potential of radium monofluoride

Abstract
The ionization potential (IP) of radium monofluoride (RaF) was measured to be 4.969(2)[10] eV, revealing a relativistic enhancement in the series of alkaline earth monofluorides. The results are in agreement with a relativistic coupled-cluster prediction of 4.969[7] eV, incorporating up to quantum electrodynamics corrections. Using the same computational methodology, an improved calculation for the dissociation energy () of 5.56[5] eV is presented. This confirms that radium monofluoride joins the small group of diatomic molecules for which , paving the way for precision control and interrogation of its Rydberg states.
The ionization potential (IP), defined as the minimum energy required to release a bound electron, is a fundamental property of atoms and molecules, and is important in chemistry and physics. The IP is central to revealing rigorous and intuitively appealing interrelationships among all electronic structural properties of atoms and molecules [1]. Electronic states of these systems can be arranged into series, which differ in their principal quantum number, , and excitation energy. The energy of states within each series increases with and eventually converges at the IP.
Due to the highly non-linear scaling of atomic and molecular properties with , states with can exhibit remarkable properties when compared to their ground states. These ‘Rydberg states’ can have micrometer-sized atomic radii, extended lifetimes and extremely large transition dipole moments, thousands of times larger than those for their low-lying (low-) states [2]. These exceptional characteristics make Rydberg states prominent systems for quantum computing and simulation [3], precision measurements [4] and the investigation of long-range interactions due to their high sensitivity to external electric fields and photon absorption.
For the majority of diatomic molecules, however, the dissociation energy () lies lower in energy than the IP. This severely limits the ability of experiments to study and manipulate their high-lying Rydberg states, as the molecules can fragment following pre-dissociation in short (ns) timescales. A notable exception to this trend, with , is BaF, which has been an important playground for the investigation of molecular Rydberg states owing to this rare property [5, 6, 7].
Molecules containing isotopes of radium have been proposed as being promising systems in which to study the fundamental symmetries of the Universe [8, 9], particularly in the hadronic sector of the Standard Model, where the rare octupole deformation of certain radium isotopes significantly boosts their sensitivity to -violating phenomena [10].
A notable example is the RaF molecule, which was theoretically predicted [8] and later experimentally confirmed to be directly laser-coolable [11, 12, 13]. Leveraging the unique properties of high-lying Rydberg states of RaF through manipulating them with external fields, could enable the sensitive control of these molecules [14, 15], offering complementary opportunities for future precision measurements. The realization of such techniques however relies upon the condition that .
Here, we demonstrate that RaF possesses an exceptionally large dissociation energy that exceeds its ionization potential. We report the first measurement of the IP of 226Ra19F, determined from the ionization threshold of its valence electron under multi-step laser excitation using both two-step and three-step ionization schemes. The obtained value is compared with ab initio calculations performed within the relativistic coupled-cluster (RCC) framework, corrected for higher-order effects, including QED contributions. The same computational method is used to calculate its dissociation energy, resulting in an improved value, thus confirming that RaF possesses the rare property in which .
Experimental details: Bunches of 226Ra19F+ were produced at the ISOLDE radioactive ion beam facility at CERN and studied with the Collinear Resonance Ionization Spectroscopy (CRIS) experiment. Details on the production of RaF molecules at ISOLDE can be found in Refs. [12, 16]. Upon entering the CRIS beamline, the RaF+ beam was neutralized in-flight after passing through a charge-exchange cell filled with a sodium vapor, where the neutral RaF molecules predominately populated the electronic ground state [17]. The non-neutralized residual ions were deflected away, while the neutral bunches entered an ultra high-vacuum interaction region ( mbar), where they were collinearly overlapped with either two or three pulsed laser beams.
The ionization threshold was measured during two separate experiments. The ionization schemes for these are shown in Fig. 1 a). In both experiments, the first laser excited the transition of the RaF molecules [12], transferring molecules residing in multiple rotational states from the vibronic ground state [11]. In the first experiment, the second laser was used to ionize molecules directly from the state, constituting a two-step resonance ionization scheme. In the second experiment, the discovery of the higher-lying state [18] allowed a three-step resonance ionization scheme to be employed. This resulted in a significantly improved signal-to-background ratio as the superior pulse energy and beam quality of the ionization laser operating in its fundamental wavelength range increased the ionization efficiency while simultaneously decreasing the non-resonant background. Additional details on the laser setups for each experiment can be found in the Supplemental Material.
The wavelengths of the ionization lasers were scanned and the resulting RaF+ molecular ions were deflected onto an ion detector. The ion count rate was monitored as a function of ionization laser wavelength. The total excitation energy delivered to each molecule by the lasers was determined by Doppler-correcting the sum of the individual photon energies of the lasers in each ionization scheme. Two scans at each ionization laser wavelength were taken; one in which all lasers of each scheme were present in the interaction region and the other one where only the ionization laser was present such that any background resulting from the ionization laser could be accounted for. The ion rate was then determined as the difference between these two scans and linearly normalized with respect to the ionization laser power.
The resulting thresholds are shown in Fig. 1 b) and c) for the two-step (teal squares) and three-step schemes (blue circles), respectively. Using the same method as presented in Ref. [19], the two sets of data were fit using a Sigmoid function and the IPs determined from the energy at which the ion rate saturates. This was done by determining the energy at which the tangent to the inflection point of the fitted Sigmoid curves intersects the maximum ion rate, and its corresponding boundaries. The IPs and their statistical uncertainties are shown as the vertical dashed-and-dotted lines and the narrow dark grey bands in Fig. 1 b) and c). A systematic uncertainty from the fitting procedure is assigned by taking the energy difference between the inflection points of the curves and the extracted IPs. This is due to the lack of consistency in the literature regarding where the IP lies with respect to measured ionization thresholds. The total combined statistical and systematic uncertainties are shown as the wide light gray bands in Fig. 1 b) and c). More details on the fitting and its justification can be found in the Supplementary Material. The ionization threshold measured using the three-step scheme (Fig. 1 c)) can be seen to increase more sharply when compared to the two-step scheme (Fig. 1 b)). This is due to the additional rotational angular momentum () selectivity introduced by having two sequential resonant excitations before the ionization step in the three-step scheme. This in turn reduced the number of high-lying Rydberg states populated by the ionization laser at energies below the IP, resulting in a sharper threshold. The ionization thresholds were determined to be 4.966(3)[10] eV and 4.970(1)[2] eV for the two-step and three-step scheme, respectively, with the statistical errors shown in curved brackets and the systematic errors shown in straight brackets.
More details on the fitting and its justification can be found in the Supplementary Material.
The majority of each molecular bunch was present in a shielded interaction region when the ionization process took place. However, a fraction of the molecules at the tail-end of each bunch were outside of this region where stray fields are non-negligible. These molecules were gated out of the data set as a conservative measure to negate any systematic downward shift of the observed thresholds resulting from the electric fields present. See the Supplementary Material for more details on the time-of-flight gating.
The potential energy curves for the , states in RaF and the RaF+ ground state calculated here have similar equilibrium bond lengths computed to be 2.246 Å, 2.191 Å and 2.167 Å, respectively. As no significant change in the molecular geometry or vibrational quantum number occurs, the observed ionization thresholds from the and state represent the adiabatic IP of RaF. The final experimental value is given as 4.969(2)[10] eV, using the weighted standard deviation of the two-step and three-step scheme measurements to calculate the statistical error. The systematic errors from the two measurements were added in quadrature as the measurements were taken from separate independent experiments under different experimental conditions.
Computational method: The relativistic single-reference coupled-cluster approach with single, double (CCSD), and perturbative triple excitations (CCSD(T)) was employed in our calculations, as implemented in the DIRAC19 program package [20]. A zero-point energy (ZPE) correction of 5 meV is added to estimate the minimum vibrational energy of RaF and RaF+ in the and states, respectively, using calculated harmonic vibrational frequencies. Relativistic core-valence-correlating Dyall basis sets [21, 22] of varying quality, cvz () augmented by a single layer of diffuse functions, were used (s-aug-cvz). The calculated potential energy curves were extrapolated to the complete basis-set limit (CBSL) using the scheme in Ref. [23] for the Dirac-Hartree-Fock energy and the scheme in Ref. [24] for the correlation contribution. In the calculations, 49 electrons were correlated and the virtual space was cut off at 50 a.u.
To correct for the limited active space used, the difference between results obtained correlating 49 electrons with a 50 a.u. cutoff and those obtained correlating all (97) electrons with a virtual space cutoff of 2000 a.u was calculated. In order to capture the full active space effect and to account for inner-core correlations, the all-electron quality basis set was used in the latter calculation. The more modestly sized dyall-cv3z and dyall-ae3z basis sets were employed, as calculations were prohibitively computationally expensive at the 4z level. Furthermore, the possible lack of diffuse functions was accounted for by taking the difference between the d-aug-cv4z and the s-aug-cv4z results. The above corrections were calculated at a single geometry point and added to the potential energy curves.
The effect of perturbative triple excitations on the CBSL level was determined to be 51 meV. The effect of full triple excitations was evaluated to be around 1 meV by comparing the IP calculated at the CCSDT and CCSD(T) levels using the MRCC code [25, 26]. These calculations were performed using the dyall.v3z basis sets, 16 correlated electrons, and a 10 a.u. virtual cutoff. Higher-level excitations were not considered owing to the very small difference between the CCSD(T) and CCSDT results.
Method | IP (eV) | (eV) |
CBS-DC-CCSD111ZPE correction is included. | 4.932 | 5.454 |
CBS-DC-CCSD(T)111ZPE correction is included. | 4.983 | 5.547 |
+aug+ae.vs.cv | 4.977 | 5.554 |
+T | 4.978 | 5.549 |
+Breit | 4.976 | 5.547 |
+QED | 4.969 | 5.556 |
Theoretical | 4.969[7] | 5.56[5] |
Experimental | 4.969(2)[10] | 5.64(24)222Scaled from Ref. [12]. |
In addition, the Breit and QED contributions were estimated. QED corrections were calculated using a development version of the DIRAC code [27]. Using this implementation, two effective QED potentials were added variationally to the DC Hamiltonian. The QED correction itself was obtained from a single-point calculation at the equilibrium geometry of neutral RaF. The Uehling potential [28] was employed for vacuum polarization and the effective potential of Flambaum and Ginges for the electron self-energy [29]. Our estimate of the size of the Breit effect relied on the fact that the electronic structure of RaF is very similar to that of Ra+ and upon ionization, the valence electron is removed from a Rydberg orbital (atomic-like and non-bonding). Thus, the contribution of the Breit interaction to the IP of RaF can be approximated by the effect calculated for the IP of Ra+ (direct calculations of molecular Breit contributions are challenging at present). These calculations were performed within the Fock-space coupled-cluster approach (DCB-FSCC), using the Tel Aviv atomic computational package [30].
The various higher-order corrections are added to the adiabatic CBS-DC-CCSD(T) IP and the final theoretical value was determined to be 4.969[7] eV as shown in Table 2. Using the same computational method, the dissociation energy of RaF was calculated to be 5.56[5] eV. The uncertainty on the calculated IP and was evaluated through further computational investigation using similar procedures that were previously employed for various properties of both atoms and molecules [31, 32, 33, 34]. The details of the uncertainty treatment for both properties can be found in the Supplemental Material.
Discussion: Fig. 2 a) shows experimental determinations of the IP of the group II monofluorides from CaF to RaF and compared with theoretical calculations using the same method as for RaF described previously [35]. Experimental IP values measured using electron-impact [36, 37, 38] and laser spectroscopy methods [39, 6] are shown as yellow diamonds and turquoise circles, respectively, with theoretical calculations shown as dark blue squares. The error bars for the laser measurements and theoretical predictions are smaller than the markers. Fig. 2 b) shows the deviation, in units of standard deviations, between the theoretical predictions (dark blue squares) and experimental values (gray line). Where they exist, the experimental values used in the comparison are from laser spectroscopy methods using Rydberg states or ionization threshold measurements, owing to their higher precision. In SrF, where no laser measurements exist, an electron-impact measurement is used for comparison. Excellent agreement within 1 standard deviation is obtained for CaF, BaF, and RaF whereas SrF agrees within 2 standard deviations. Fig. 2 a) highlights the ability to determine the IP of RaF at a comparable or better precision than experiments on the homologues, despite the technical challenges imposed by its short-lived nature.
The IP of the group II monofluorides decreases progressively with increasing proton number, before increasing again in RaF. This enhancement in valence electron binding energy can be attributed to relativistic effects, which are significantly more pronounced in the heavier RaF. These effects result from the relativistic increase of the electron mass driven by the high velocity that bound electrons experience in atoms and molecules containing heavy elements [40]. This in turn spatially contracts orbitals with low angular momentum which increases their binding energies [41, 42]. The relativistic treatment required for correctly describing RaF is fully achieved in this study, yielding good agreement between theory and experiment as in the case of its homologues.
Fig. 2 c) shows how the calculated IP for RaF evolves with an increasing level of computational sophistication. An agreement within 1 standard deviation is found between experiment and CCSD(T) theory that incorporates a complete basis set-extrapolation correction, non-perturbative triple excitations, and Breit and QED corrections. The final theoretical uncertainty is 7 meV, owing to the extensive implementation of higher-order contributions.
Benchmarking the predictive power of molecular theory is crucial for the study of radioactive molecules, which is a prominent future avenue for precision tests of the Standard Model [43]. As most of the short-lived radioactive molecules of interest can only be made in small quantities at specialized facilities, none of their chemical properties are known experimentally even though knowledge of their chemical behavior is necessary to guide, or even enable altogether, their production and study [44, 45]. Therefore, while the calculated IP of RaF is accurate even for predictions of lower sophistication in Fig. 2, the ability to reduce the theoretical uncertainty by including higher-order contributions, up to the QED level, remains important to aid experiments attempting to perform the first spectroscopy of previously unexplored systems.
The previously calculated values of the dissociation energy, , for RaF vary between 3.980 eV [8] and 5.282 eV [46] with no corresponding uncertainties reported. Using the same computational method as for the IP, which achieves a more extensive and rigorous treatment of higher-order contributions compared to previous studies, an improved calculation of the RaF equal to 5.56[5] eV is presented. This is in agreement with the RaF extracted from Morse potential fits in Ref. [12], provided they are scaled by 1.55(6). Dissociation energy determinations from Morse potential fits use linear extrapolations that significantly underestimate the true dissociation energy in molecules where ionic bonding is important [47, 48]. By taking the extrapolated values for CaF [49], SrF [50] and BaF [51], and comparing them to the true measured dissociation energies [52], the aforementioned scaling factor is obtained, which results in a RaF of 5.64(24) eV (using the value from Ref. [12]). The resulting value and uncertainty of the computed , corroborated by the agreement between experiment and theory for the IP, confirms that RaF joins BaF in the rare class of diatomic molecules for which IP.
Conclusion: The ionization potential of radium monofluoride (RaF) was measured with the CRIS experiment at CERN-ISOLDE to be 4.969(2)[10] eV, through consistent observations of the ionization thresholds starting from two different excited electronic states. The experimental results were compared against predictions calculated with single-reference relativistic coupled-cluster theory that incorporates triple excitations, basis-set, Breit-effect and QED corrections. A good agreement is obtained between experiment and theory for RaF, in addition to its lighter homologues. The dissociation energy was found to lie above the IP, offering the possibility of accessing and utilizing Rydberg states in RaF, without loss due to pre-dissociation. This opens up several opportunities for controlling and manipulating the molecule via external fields [14, 15]. Additionally, these states can also offer an alternative window into probing the structure of radium nuclei [5].
Acknowledgments: The authors thank Michael Morse for his insights. This work was supported by the Office of Nuclear Physics, U.S. Department of Energy, under grants DE-SC0021176 and DE-SC0021179, the MISTI Global Seed Funds, Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Projektnummer 328961117 – SFB 1319 ELCH; STFC grants ST/P004423/1 and ST/V001116/1, ERC Consolidator Grant no. 648381 (FNPMLS), Belgian Excellence of Science (EOS) project No. 40007501; KU Leuven C1 project No. C14/22/104, FWO project No. G081422N, International Research Infrastructures (IRI) project No. I001323N, the European Unions Grant Agreement 654002 (ENSAR2), LISA: European Union’s H2020 Framework Programme under grant agreement no. 861198, The Swedish Research Council (2016-03650 and 2020-03505), The National Key RD Program of China (No: 2022YFA1604800) (X.F.Y.) and the National Natural Science Foundation of China (No:12027809), the Dutch Research Council (NWO) projects number VI.C.212.016 and Vi.Vidi.192.088, the Slovak Research and Development Agency projects APVV-20-0098 and APVV-20-0127, JSPS Overseas Challenge Program for Young Researchers, Grant No. 201880193, JST Moonshot R&D Grant No. JPMJMS2269 and NSF RII Track-4 (Grant No. 2327247). We thank the Center for Information Technology at the University of Groningen for their support and for providing access to the Peregrine and Hábrók high performance computing cluster.

### v3

@@
 The results are in agreement with a relativistic coupled-cluster prediction of 4.969[7] eV, incorporating up to quantum electrodynamics corrections.
-Using the same computational methodology, an improved calculation for the dissociation energy () of 5.56[5] eV is presented.
+Using the same computational methodology, an improved calculation for the dissociation energy () of 5.54[5] eV is presented.
 This confirms that radium monofluoride joins the small group of diatomic molecules for which , paving the way for precision control and interrogation of its Rydberg states.
@@
 A notable exception to this trend, with , is BaF, which has been an important playground for the investigation of molecular Rydberg states owing to this rare property [5, 6, 7].
-Molecules containing isotopes of radium have been proposed as being promising systems in which to study the fundamental symmetries of the Universe [8, 9], particularly in the hadronic sector of the Standard Model, where the rare octupole deformation of certain radium isotopes significantly boosts their sensitivity to -violating phenomena [10].
+Molecules containing isotopes of radium have been proposed as being promising systems in which to study the fundamental symmetries of the Universe [8, 9, Wilkins2023], particularly in the hadronic sector of the Standard Model, where the rare octupole deformation of certain radium isotopes significantly boosts their sensitivity to -violating phenomena [10].
 A notable example is the RaF molecule, which was theoretically predicted [8] and later experimentally confirmed to be directly laser-coolable [11, 12, 13].
@@
 Method | IP (eV) | (eV) |
-CBS-DC-CCSD111ZPE correction is included. | 4.932 | 5.454 |
-CBS-DC-CCSD(T)111ZPE correction is included. | 4.983 | 5.547 |
-+aug+ae.vs.cv | 4.977 | 5.554 |
-+T | 4.978 | 5.549 |
-+Breit | 4.976 | 5.547 |
-+QED | 4.969 | 5.556 |
-Theoretical | 4.969[7] | 5.56[5] |
-Experimental | 4.969(2)[10] | 5.64(24)222Scaled from Ref. [12]. |
+CBS-DC-CCSD | 4.932111ZPE correction is included. | 5.427 |
+CBS-DC-CCSD(T) | 4.983111ZPE correction is included. | 5.520 |
++aug+ae.vs.cv | 4.977 | 5.538 |
++T | 4.978 | 5.534 |
++Breit | 4.976 | 5.532 |
++QED | 4.969 | 5.541 |
+Theoretical | 4.969[7] | 5.54[5] |
+Experimental | 4.969(2)[10] | 5.61(24)222Scaled from Ref. [12]. |
 In addition, the Breit and QED contributions were estimated.
@@
 The various higher-order corrections are added to the adiabatic CBS-DC-CCSD(T) IP and the final theoretical value was determined to be 4.969[7] eV as shown in Table 2.
-Using the same computational method, the dissociation energy of RaF was calculated to be 5.56[5] eV.
+Using the same computational method, the dissociation energy of RaF was calculated to be 5.54[5] eV.
 The uncertainty on the calculated IP and was evaluated through further computational investigation using similar procedures that were previously employed for various properties of both atoms and molecules [31, 32, 33, 34].
@@
 2, the ability to reduce the theoretical uncertainty by including higher-order contributions, up to the QED level, remains important to aid experiments attempting to perform the first spectroscopy of previously unexplored systems.
-The previously calculated values of the dissociation energy, , for RaF vary between 3.980 eV [8] and 5.282 eV [46] with no corresponding uncertainties reported.
-Using the same computational method as for the IP, which achieves a more extensive and rigorous treatment of higher-order contributions compared to previous studies, an improved calculation of the RaF equal to 5.56[5] eV is presented.
+The previously calculated values of the ZPE-corrected dissociation energy, , for RaF vary between 3.954 eV [8] and 5.255 eV [17] with no corresponding uncertainties reported.
+Using the same computational method as for the IP, which achieves a more extensive and rigorous treatment of higher-order contributions compared to previous studies, an improved calculation of the RaF equal to 5.54[5] eV is presented.
 This is in agreement with the RaF extracted from Morse potential fits in Ref.
@@
 Dissociation energy determinations from Morse potential fits use linear extrapolations that significantly underestimate the true dissociation energy in molecules where ionic bonding is important [46, 47].
-By taking the extrapolated values for CaF [49], SrF [50] and BaF [51], and comparing them to the true measured dissociation energies [52], the aforementioned scaling factor is obtained, which results in a RaF of 5.64(24) eV (using the value from Ref.
+By taking the extrapolated values for CaF [48], SrF [49] and BaF [50], and comparing them to the true measured dissociation energies [51], the aforementioned scaling factor is obtained, which results in a RaF of 5.61(24) eV (using the value from Ref.
 [12]).
@@
 Conclusion: The ionization potential of radium monofluoride (RaF) was measured with the CRIS experiment at CERN-ISOLDE to be 4.969(2)[10] eV, through consistent observations of the ionization thresholds starting from two different excited electronic states.
-The experimental results were compared against predictions calculated with single-reference relativistic coupled-cluster theory that incorporates triple excitations, basis-set, Breit-effect and QED corrections.
-A good agreement is obtained between experiment and theory for RaF, in addition to its lighter homologues.
 The dissociation energy was found to lie above the IP, offering the possibility of accessing and utilizing Rydberg states in RaF, without loss due to pre-dissociation.